OpenStudy (anonymous):

Factorise Fully (sin x)^3-2.25*sin x

OpenStudy (cwrw238):

first take out sinx

OpenStudy (cwrw238):

= sin x(sin^2x - 2.25)

OpenStudy (cwrw238):

something wrong with openstudy..

OpenStudy (anonymous):

yes what else ???

OpenStudy (cwrw238):

try again - the part in the parenthesis is difference of 2 squares - use the identity a^2 - b^2 = (a + b)(a - b) - let a = sin x and b = 1.5

OpenStudy (cwrw238):

- I can only write 2 lines - openstudy wont allow more!

OpenStudy (cwrw238):

1.5 because its the square root of 2.25

OpenStudy (cwrw238):

- just replace a with sin x and b with 1.5 in the formula I posted

OpenStudy (cwrw238):